# A Comprehensive Graph Pooling Benchmark: Effectiveness, Robustness and Generalizability

### v1

@@ 1 Introduction
Recently, graph neural networks (GNNs) have garnered significant attention due to their remarkable ability to process graph-structured data across various domains [1, 2] including social networks [3, 4, 5], rumor detection [6, 7], biological networks [8, 9], recommender systems [10] and community detection [11, 12]. Graph pooling approaches play a crucial role in GNNs by enabling the hierarchical reduction of graph representations, which is essential for capturing multi-scale structures and long-range dependencies [13, 14, 15]. They can preserve crucial topological semantics and relationships, which have shown effective for tasks including graph classification, node clustering, and graph generation [13, 16, 17, 2]. In addition, by aggregating nodes and edges, graph pooling can also simplify large-scale graphs, facilitating the application of GNNs in real-world problems [18, 19, 20, 21, 22]. Therefore, understanding and enhancing graph pooling approaches is the key to increasing GNN performance across various domains, driving the progress of deep geometric learning.
@@
Towards this end, we present a comprehensive graph pooling benchmark, which includes 15 graph pooling methods and 21 datasets across different graph machine learning problems. In particular, we extensively investigate graph pooling approaches across three key dimensions, i.e., effectiveness, robustness, and generalizability. To begin, we provide a fair and thorough effectiveness comparison of existing graph pooling approaches across graph classification, graph regression and node classification. Then, we evaluate the robustness of graph pooling approaches under both noise attacks on graph structures and node attributes. In addition, we study the generalizability of different approaches under out-of-distribution shifts from both size and density levels. Finally, we include efficiency comparison, parameter analysis and backbone analysis for completeness.
From extensive experiments, we have four observations as follows: (1) Dense pooling approaches generally outperform pooling on graph classification and graph regression. (2) Different graph pooling approaches have a limited impact on node classification. (3) Feature masking influences the performance of graph pooling more seriously compared with structure perturbation and dense pooling is more robust to noise attack. (4) Most of graph pooling approaches suffer from serious performance decrement from distribution shifts.
@@ 4.2 Effectiveness Analysis
Performance on Graph Classification. To begin, we investigate the performance of different graph pooling approaches on graph classification. The results of compared approaches on seven popular datasets are recorded in Table 2. From the results, we have the following observations. Firstly, in general, dense pooling methods outperform sparse pooling methods. However, dense pooling methods encounter memory bottlenecks when processing large datasets with tens of thousands of graphs, such as COLORS-3. Secondly, two dense pooling approaches, AsymCheegerCutPool and DiffPool are among the top out of all the compared methods, achieving optimal or near-optimal results on five datasets. The potential reasons include AsymCheegerCutPool minimizing the graph total variation of the cluster assignments, and DiffPool improving the quality of membership assignments by using a link prediction auxiliary objective [33, 32]. Thirdly, among the sparse pooling methods, ASAPool, CGIPool, and HGPSLPool demonstrate the best overall performance. CGIPool achieves the best results on IMDB-M, while HGPSLPool achieves the best results on BZR.
@@
Performance on Graph Regression. We further explore the performance of different pooling methods through graph-level regression tasks. As shown in Table 3, we can observe the following: Firstly, dense pooling methods significantly outperform sparse pooling methods. DMoNPool, HoscPool, and JustBalancePool are the best-performing pooling methods, with DMoNPool achieving optimal or near-optimal results on four datasets, and JustBalancePool achieving optimal or near-optimal results on three datasets. Secondly, KMISPool performs well among sparse pooling methods, surpassing AsymCheegerCutPool in dense pooling methods.
Performance on Node Classification. Table 4 presents the performance of various sparse pooling methods in node classification tasks. We can observe that: Firstly, different pooling approaches have consistent performance across these datasets. The potential reason is that graph convolution is the core of semantics exploration on the node level. Secondly, TopKPool demonstrates the best performance among the methods used, potentially because it can adapt to the irregularity of graph data and extract information from the k most important nodes [23]. GSAPool achieves the best or second-best accuracy in 5 out of 7 datasets, as previously mentioned studies have indicated. Before discarding less important nodes, GSAPool aggregates the features of nodes, ensuring that the pooled nodes contain sufficient effective graph information [30]. This aggregation method ensures that the selected nodes not only contain their own features but also integrate information from neighboring nodes. Thirdly, the scalability of ASAPool, PANPool, and HGPSLPool still requires improvement, as they are unable to complete training on larger datasets such as GitHub.
@@ 4.3 Robustness Analysis
The compared performance for three types of random noise on eight graph pooling methods on the PROTEINS, NCI1, NCI109, ENZYMES, and BZR datasets are shown in Table LABEL:noise. With a probability of 50%, edges of the graph are randomly removed or added, and node features are randomly masked with the same likelihood. From the results, we have the following observations. Firstly, masking node features would degrade the performance the most, which demonstrates that node attributes are still the most crucial for graph representation learning. Secondly, the performance ranking of different methods is consistent under different noise attacks but not consistent across different datasets. This indicates that we can choose the optimal graph pooling approaches by validating the performance under any type of noise attack in practice. Thirdly, overall, dense pooling methods exhibit stronger resistance to random noise attacks compared to sparse pooling methods. For instance, JustBalancePool achieves optimal or near-optimal results on three datasets, and DiffPool also performs exceptionally well on most datasets. Notably, sparse pooling methods significantly outperform dense pooling methods on the multi-classification dataset ENZYMES, as evidenced by the performance of TopKPool and COPool. This suggests that the choice of pooling method can be tailored based on the nature of the classification task.
As depicted in Figure 3, the model’s performance generally declines as the noise intensity increases. It is observed that at the same level of noise, the impact on accuracy is more pronounced on NCI109, while it is relatively minor on ENZYMES. The accuracy degradation due to edge addition noise is less than that caused by node masking noise and edge deletion noise. Among the three types of noise, although the accuracy of nearly all methods decreases amidst fluctuations, JustbalancePool and DiffPool exhibit the strongest robustness. For the NCI1 dataset, MincutPool shows significant variance in accuracy performance.
@@ 4.4 Generalizability Analysis
Table 6 presents the performance of all 15 graph pooling methods under out-of-distribution shifts. For the datasets D&D, PROTEINS, NCI1, and NCI109, two types of distribution shifts are implemented. The first type is based on the number of nodes, where the smallest 50% of graphs by node count are used as the training set, and the largest 20% as the test set, with the remainder serving as the validation set [64, 65]. Following the same criteria, the second type of out-of-distribution shifts are generated based on graph density [65]. For further details, please refer to the Appendix.
From Table 6, we have the following observations. Firstly, the majority of graph pooling approaches suffer from serious performance decreases when it comes to out-of-distribution shifts. For example, most of the approaches have the accuracy below 35% when it comes to size-level distribution shifts. Secondly, for the size-based out-of-distribution shift, KMISPool exhibits superior performance. The underlying reason is its theoretical guarantee of path length distortion boundaries, as well as its ability to preserve key topological features in the coarsened graph. This endows it with enhanced performance when facing size-based out-of-distribution shifts [29]. Thirdly, for the density-based out-of-distribution shift, DiffPool demonstrates optimal performance. The potential reason is its stronger capability to preserve the information of the original graph [33, 41].
@@ 4.5 Further Analysis
Efficiency Comparison. In this part, we conduct an efficiency analysis of graph pooling methods on the ENZYMES and NCI1 datasets. We calculate the time of the algorithms by measuring the duration needed to complete 200 epochs of training with the 256 batch size. For space efficiency, we compute the GPU memory utilization during the training process. As shown in Figure 4, dense pooling methods such as DiffPool, MincutPool, and JustBalancePool have significantly higher time and space costs. Notably, although KMISPool performs poorly on the multi-classification dataset ENZYMES, it has one of the lowest time and space costs across all three datasets. Conversely, COPool has the highest GPU memory usage among the eight selected methods.
@@ A.1 Graph Classification and Graph Regression
The characteristics of a graph are typically described by the properties of its substructures and the interactions between them [70]. Unlike node classification tasks, graph classification or graph regression tasks require attention to the global information of the graph [41, 39]. This necessitates a graph pooling mechanism capable of extracting such global information [38]. Currently, extensive research has been conducted on applying graph pooling techniques to graph classification and regression tasks. These methods can be broadly categorized into two main streams: flat pooling and hierarchical pooling [38, 39].
Flat pooling, as the name suggests, generally involves globally aggregating the features of all nodes in the graph without considering the graph’s structure [20]. It directly combines information from all nodes into a single global vector representation. The advantage of this approach lies in its simplicity and efficiency, which is often sufficiently effective for many application scenarios [39]. Classical flat pooling methods include Set2set, SortPool, DeepSet, and DAGCN.
Set2set’s strength is its use of the chain rule to effectively represent the joint probability of sequences, naturally accommodating variable-sized inputs and/or outputs in sequence form. This method is particularly suitable for scenarios requiring the handling of variable-length inputs [71]. SortPool’s advantage is that it sorts graph vertices in a consistent order, enabling the training of traditional neural networks on graphs. This sorting mechanism allows SortPool to better preserve graph structure information [72]. DeepSet frames the problem within the setting of set learning, encompassing a simple yet powerful readout layer formula. This formula can encode or approximate any continuous permutation-invariant function over a set [73]. DAGCN uses an attention-based graph convolution layer to automatically learn the importance of neighbors at different hops, followed by a self-attention pooling layer that generalizes graph representations from various aspects of the matrix graph embedding, thereby retaining as much of each node’s features and inter-node topological structure as possible [74].
In addition to these classic methods, some more novel flat pooling approaches have been proposed. For instance, GMT is a global pooling layer based on multi-head attention that captures node interactions according to structural dependencies between nodes, while satisfying injectivity and permutation invariance [75]. It also boasts high memory utilization efficiency. QSGCNN extracts multi-scale vertex features from the perspective of quantum information propagation between grid vertices in each graph, integrating graph representation and learning in quantum space graph convolution layers, which do not alter the original spatial positions of vertices [76]. GraphTrans employs Transformer-based self-attention to learn long-range pairwise relationships, effectively addressing optimization instability [77]. DMLAP adapts to both local and global structural information in the graph, featuring an attention pooling layer for each message passing step and computing the final graph representation through unified layer-by-layer graph representationy [78].
Hierarchical pooling, in contrast to flat pooling, incrementally simplifies graph information layer by layer [79, 41]. For a three-layer graph convolutional network (GCN), there are typically two hierarchical pooling layers interleaved between the three convolutional layers. This gradual simplification allows hierarchical pooling to better preserve the hierarchical structure and local information of the graph, thus extracting meaningful global representations at higher levels [79, 38]. This method has significant advantages when dealing with complex graph structures, as it can better capture the intricate relationships between nodes and substructures through multi-level simplification and aggregation [33, 39].
Hierarchical pooling methods for graph classification and regression have been extensively developed, with notable methods including TopKPool, EdgePool, SAGPool, AttPool, ASAPool, PANPool, MVPool, and LiftPool. TopKPool, as one of the most classical hierarchical pooling methods, selects the most important nodes by learning projection scores, thereby retaining crucial information while reducing the graph’s size. This method is applicable to various graph structures, including large-scale and sparse graphs [23]. EdgePool learns local and sparse pooling transformations, which can be integrated into existing architectures without necessitating any changes to the training process [80]. SAGPool is a self-attention-based graph pooling method that leverages the self-attention mechanism of graph convolution, allowing the pooling method to simultaneously consider node features and graph topology [24]. AttPool adaptively selects nodes significant to graph representation and generates hierarchical features by aggregating attention-weighted information from nodes, excelling in learning hierarchical representations of graph embeddings [81]. ASAPool utilizes a novel self-attention network and an improved GNN formula to capture the importance of each node within the given graph. It also learns sparse soft clustering assignments for each layer’s nodes, effectively pooling subgraphs to form a pooled graph [25]. PANPool extends the Laplacian graph to a new transition matrix, which can be customized for different graph data with varying sizes and structures [26]. MVPool is based on a multi-view graph pooling operator that uses attention mechanisms to facilitate cooperation among different perspectives to produce robust node rankings. The pooling operation then adaptively selects a subset of nodes to form an induced subgraph based on the ranking list [82]. LiftPool enhances hierarchical graph representations by maximizing the retention of local structural information within the graph pool. It introduces an additional graph lifting stage before graph coarsening to preserve the local information of the removed nodes and decouples node removal from feature reduction processes [83].
Despite the ongoing innovations in graph pooling for graph classification and regression, several challenges persist. Firstly, the time complexity and space complexity of algorithms remain critical issues [41]. Secondly, most methods are designed for standard graphs, whereas real-world datasets include many other types of graphs with different characteristics and structures [39]. Thirdly, the interpretability of graph pooling requires improvement. Currently, many graph pooling methods cannot effectively separate noise information from the input graph, and their performance lacks robustness when facing attacks on graph topology or features [38].
In conclusion, while hierarchical pooling methods have significantly advanced graph classification and regression tasks, addressing these challenges is crucial for further progress. Improving the efficiency, applicability, and interpretability of these methods will enhance their utility in diverse and complex real-world scenarios.
@@ A.2 Node Classification
Graph node classification aims to classify individual nodes within a graph by predicting their labels based on node attributes and their relationships [84]. For example, in a social network, one might aim to predict each user’s political inclination, while in a protein-protein interaction network, the goal might be to predict each protein’s functional role [85].
Flat pooling methods are less commonly applied to node classification tasks. Among those that do exist, FusionPool is a representative method. FusionPool is designed to address the high parameter count and computational complexity inherent in higher-order graph convolutional networks (GCNs). It achieves this by non-linearly integrating information from both low-order and high-order GCNs [86]. FusionPool combines neighborhood information matrices from different orders, effectively balancing the complexity and richness of the learned node representations.
Hierarchical Pooling Methods for Node Classification Several hierarchical pooling methods have been specifically designed for node classification [41]. These methods aim to capture and preserve the graph’s local and global structural information while simplifying the graph. Representative methods include SEPool, GRAHIES, MVPool, VIPool, TopKPool, DHT, and EdgeCut. SEPool addresses the local structure damage and suboptimal issues inherent in hierarchical pooling methods. It employs the concept of structural entropy and designs a global optimization algorithm that generates a clustering assignment matrix for pooling in a single step, without the need for layer-specific compression ratios [87]. GRAHIES captures the inherent hierarchical topological features of many real-world graphs by merging hierarchical node embeddings. It adaptively learns the multi-level hierarchical structure of the input graph and combines graph representations from different hierarchical levels to capture the intrinsic global hierarchical structure of the original graph [88]. VIPool creates multi-scale graphs in a trainable manner and introduces a novel feature crossing layer that enables cross-scale feature exchange. It selects the most informative subset of vertices based on a neural estimation of mutual information between vertex features and their neighborhoods [89]. VIPool integrates intermediate features between two scales to achieve mutual enhancement. DHT transforms the edges of a graph into nodes of a hypergraph. This dual hypergraph structure allows the application of node representation message-passing techniques to edges. After obtaining edge representations from the hypergraph, it clusters or drops edges to derive an overall graph-level edge representation [90]. EdgeCut generates different versions of the input graph at various scales. It computes edge scores corresponding to the importance of edges during GNN information propagation, addressing the issue of nodes having unordered and irregular neighborhoods [91].
@@ Appendix B Details of Selected Pooling Methods
TopKPool Firstly, a score is computed for each node. These scores can be obtained through a combination of node features and a learnable parameter vector. Based on the computed scores, the top k nodes with the highest scores are selected. Once the top k nodes are determined, the subgraph consisting of these nodes and their associated edges is extracted [23]. This subgraph preserves the local structure of the original graph while significantly reducing the number of nodes. The primary purpose of this step is to simplify the graph structure, making subsequent computations more efficient, and to embed the subgraph into a lower-dimensional space [23].
SAGPool Firstly, a self-attention mechanism is employed to assess the importance of nodes. Based on the importance scores of the nodes, a subset of nodes is selected for pooling. Typically, the top k nodes with the highest importance scores are chosen, and the selected nodes along with their corresponding edges form a new subgraph [24]. This subgraph preserves the critical structural information of the original graph while reducing its size. Finally, feature aggregation is performed on the selected nodes to obtain the final output [24].
ASAPool The ASAPool aims to enhance the aggregation effect of feature maps by introducing a spatial attention mechanism. This selectively focuses on important features, thereby strengthening the model’s expressive power. For the input features, convolutional operations are applied to the feature maps to extract local features [25]. The convolutional kernel size is typically the same as the pooling window size. Then, a lightweight attention module is used to compute the attention weights for each spatial position. This can be implemented using convolutional layers or fully-connected layers to generate the weight matrix [25]. Finally, the attention weight matrix is applied to the input feature map, yielding the weighted feature map [25].
PANPool includes a convolution operation that involves paths linking each message sender and receiver, with these paths having learnable weights depending on their length, corresponding to a maximum entropy random walk. It extends the graph Laplacian to what is referred to as the Maximum Entropy Transition (MET) matrix, whose diagonal entries are directly related to subgraph centrality, thereby providing a natural adaptive pooling mechanism [26].
COPool combines pooled representations learned from both node and edge views. Through cross-view interactions, the edge view pooling and node view pooling mutually enhance each other to learn more informative graph-level representations [27].
CGIPool combines pooled representations learned from both node and edge views. Through cross-view interactions, the edge view pooling and node view pooling mutually enhance each other to learn more informative graph-level representations [28].
KMISPool enhances the reduction sampling mechanism by applying downsampling to graph data. This graph pooling method corresponds to a controllable isometric coarsening mechanism in regular data structures [29].
GSAPool simultaneously considers both the structural and feature information of the graph. It aggregates node feature information before discarding unimportant nodes; thus, the selected nodes contain information from neighboring nodes, which can enhance the utilization of features from the unselected nodes [30].
HGPSLPool integrates graph pooling and structure learning into a unified module to generate hierarchical representations of the graph. Through the graph pooling operation, it adaptively selects a subset of nodes to form the induced subgraph for subsequent layers [31].
AsymCheegerCutPool computes cluster assignments by optimizing a tighter relaxation of the minimum cut based on Graph Total Variation (GTV). These cluster assignments can be directly used for vertex clustering or to implement graph pooling [32].
DiffPool is a differentiable graph pooling module capable of generating hierarchical representations of a graph. It can be integrated end-to-end with various graph neural network architectures. DiffPool learns differentiable soft cluster assignments for nodes at each layer of a deep GNN, mapping nodes to a set of clusters, which then form the coarsened input for the next GNN layer [33].
MincutPool formulates a continuous relaxation of the normalized minCUT problem and trains a GNN to compute cluster assignments that minimize this objective. It does not require spectral decomposition and learns a clustering function that can be quickly evaluated on out-of-sample graphs [34].
DMoNPool is an unsupervised pooling method inspired by the modularity measure of clustering quality, which is typically used to recover cluster structures [35].
HoscPool is a cluster-based graph pooling operator that hierarchically captures higher-order information, resulting in richer graph representations. It end-to-end learns a probabilistic clustering assignment matrix by minimizing a relaxed formulation of motif spectral clustering in the objective function, which is then extended to the pooling operator [36].
JustBalancePool optimizes an unsupervised loss composed of two terms. The first term ensures that connected nodes are assigned to the same cluster, while the second is a balancing term that prevents degenerate solutions by encouraging samples to be assigned to only one cluster and clusters to have similar sizes [37].

### v2

@@ Abstract
Keywords Graph Pooling; Benchmark; Graph Neural Networks; Graph Machine Learning
@@ 1 Introduction
Recently, graph neural networks (GNNs) have garnered significant attention with extensive benchmarks [1, 2, 3] due to their remarkable ability to process graph-structured data across various domains including social networks [4, 5, 6, 7], rumor detection [8, 9], biological networks [10, 11], recommender systems [12] and community detection [13, 14]. Graph pooling approaches play a crucial role in GNNs by enabling the hierarchical reduction of graph representations, which is essential for capturing multi-scale structures and long-range dependencies [15, 16, 17]. They can preserve crucial topological semantics and relationships, which have shown effective for tasks including graph classification, node clustering, and graph generation [15, 18, 19, 2]. In addition, by aggregating nodes and edges, graph pooling can also simplify large-scale graphs, facilitating the application of GNNs in real-world problems [20, 21, 22, 23, 24]. Therefore, understanding and enhancing graph pooling approaches is the key to increasing GNN performance, driving the progress of deep geometric learning.
@@
Towards this end, we present a comprehensive graph pooling benchmark, which includes 15 graph pooling methods and 21 datasets across different graph machine learning problems. In particular, we extensively investigate graph pooling approaches across three key perspectives, i.e., effectiveness, robustness, and generalizability. To begin, we provide a fair and thorough effectiveness comparison of existing graph pooling approaches across graph classification, graph regression and node classification. Then, we evaluate the robustness of graph pooling approaches under both noise attacks on graph structures and node attributes. In addition, we study the generalizability of different approaches under out-of-distribution shifts from both size and density levels. Finally, we include efficiency comparison, parameter analysis and backbone analysis for completeness.
From extensive experimental results, we have four observations as follows: (1) Dense pooling approaches generally outperform pooling on graph classification and graph regression. (2) Different graph pooling approaches have a limited impact on node classification. (3) Feature masking influences the performance of graph pooling more seriously compared with structure perturbation and dense pooling is more robust to noise attack. (4) Most of graph pooling approaches suffer from serious performance decrement from distribution shifts.
@@ 4.2 Effectiveness Analysis
Performance on Graph Classification. To begin, we investigate the performance of different graph pooling approaches on graph classification. The results of compared approaches on seven popular datasets are recorded in Table 2. From the results, we have the following observations. Firstly, in general, dense pooling methods outperform sparse pooling methods. However, dense pooling methods encounter memory bottlenecks when processing large datasets with tens of thousands of graphs, such as COLORS-3. Secondly, two dense pooling approaches, AsymCheegerCutPool and DiffPool are among the top out of all the compared methods, achieving optimal or near-optimal results on five datasets. The potential reasons include that AsymCheegerCutPool minimizes the graph total variation of the cluster assignments, and DiffPool improves the quality of membership assignments by using a link prediction auxiliary objective [35, 34]. Thirdly, among the sparse pooling methods, ASAPool, CGIPool, and HGPSLPool demonstrate the best overall performance. To be specific, CGIPool achieves the best results on IMDB-M, while HGPSLPool achieves the best results on BZR.
@@
Performance on Graph Regression. We further explore the performance of different pooling methods through graph-level regression tasks. As shown in Table 3, we can observe the following: Firstly, dense pooling methods significantly outperform sparse pooling methods. DMoNPool, HoscPool, and JustBalancePool are the best-performing pooling methods, with DMoNPool achieving optimal or near-optimal results on four datasets, and JustBalancePool achieving optimal or near-optimal results on three datasets. Secondly, KMISPool performs well among sparse pooling methods, outperforming AsymCheegerCutPool in dense pooling methods.
Performance on Node Classification. Table 4 presents the performance of various sparse pooling methods in node classification tasks. We can observe that: Firstly, different pooling approaches have consistent performance across these datasets. The potential reason is that graph convolution is the core of semantics exploration on the node level. Secondly, TopKPool demonstrates the best performance among the methods used, potentially because it can adapt to the irregularity of graph data and extract information from the k most important nodes [25]. GSAPool achieves the best or second-best accuracy in 5 out of 7 datasets, as previously mentioned studies have indicated. Before discarding less important nodes, GSAPool aggregates the features of nodes, ensuring that the pooled nodes contain sufficient effective graph information [32]. This aggregation strategy ensures that the selected nodes not only contain their own features but also integrate information from neighboring nodes. Thirdly, the scalability of ASAPool, PANPool, and HGPSLPool still requires improvement, as they are unable to complete training on larger datasets such as GitHub.
@@ 4.3 Robustness Analysis
The compared performance for three types of random noise on eight graph pooling methods on the PROTEINS, NCI1, NCI109, ENZYMES, and BZR datasets are shown in Table LABEL:noise. With a probability of 50%, edges of the graph are randomly removed or added, and node features are randomly masked with the same likelihood. From the results, we have the following observations. Firstly, masking node features would degrade the performance the most, which demonstrates that node attributes are still the most crucial for graph representation learning. Secondly, the performance ranking of different methods is consistent under different noise attacks but not consistent across different datasets. This indicates that we can choose the optimal graph pooling approaches by validating the performance under any type of noise attack in practice. Thirdly, overall, dense pooling methods exhibit stronger resistance to random noise attacks compared to sparse pooling methods. For instance, JustBalancePool achieves optimal or near-optimal results on three datasets, and DiffPool also performs exceptionally well on most datasets. In addition, sparse pooling methods significantly outperform dense pooling methods on the multi-classification dataset ENZYMES, as evidenced by the performance of TopKPool and COPool. This indicate that the choice of pooling method can be tailored based on different classification tasks.
As depicted in Figure 3, the model’s performance generally declines as the noise intensity increases. It is observed that at the same level of noise, the accuracy decrease is more evident on NCI109, while it is relatively minor on ENZYMES. The accuracy degradation due to edge addition noise is less than that caused by node masking noise and edge deletion noise. Among the three types of noise, although the accuracy of nearly all methods decreases with fluctuations, JustbalancePool and DiffPool exhibit the strongest robustness. For the NCI1 dataset, MincutPool shows significant variance in terms of accuracy.
@@ 4.4 Generalizability Analysis
Table 6 presents the performance of all 15 graph pooling methods under out-of-distribution shifts. For the datasets D&D, PROTEINS, NCI1, and NCI109, two types of distribution shifts are implemented. The first type is based on the number of nodes, where the smallest 50% of graphs by node count are used as the training set, and the largest 20% as the test set, with the remaining graphs serving as the validation set [66, 67]. Following the same criteria, the second type of out-of-distribution shifts are generated based on graph density [67]. For further details, please refer to the Appendix.
From Table 6, we have the following observations. Firstly, the majority of graph pooling approaches suffer from serious performance decreases when it comes to out-of-distribution shifts. For example, most of the approaches have the accuracy below 35% when it comes to size-level distribution shifts. Secondly, for the size-based out-of-distribution shift, KMISPool exhibits superior performance compared with the other approaches in all cases. The underlying reason is its ability to preserve key topological features in the coarsened graph [31]. Thirdly, for the density-based out-of-distribution shift, DiffPool demonstrates optimal performance. The potential reason is that it adopts a learnable graph clustering process to adaptively explore hierarchical semantics [35, 43].
@@ 4.5 Further Analysis
Efficiency Comparison. In this part, we conduct an efficiency analysis of graph pooling methods on the ENZYMES and NCI1 datasets. We calculate the time of different algorithms needed to complete 200 epochs of training with the 256 batch size. For space efficiency, we compute the GPU memory utilization during the training process. The compared results are shown in Figure 4. From the results, dense pooling methods such as DiffPool, MincutPool, and JustBalancePool have significantly higher time and space costs. Notably, although KMISPool performs poorly on the multi-classification dataset ENZYMES, it has one of the lowest time and space costs across all three datasets. Conversely, COPool has the highest GPU memory usage among the eight compared methods.
@@ A.1 Graph Classification and Graph Regression
Graphs provide an effective tool to represent interaction among different objects [72]. Graph classification [73] is a fundamental graph machine learning problem, which aims to classify each graph into its corresponding category. The majority of current works adopt message passing mechanisms [72], where each node receives information from its neighbors in a recursive manner. Then, a graph readout function is adopted to summarize all node representations into a graph-level representation for downstream classification. Graph classification has extensive applications in various domains such as molecular property prediction [74] and protein function analysis [75]. Graph regression [76] is close to graph classification which maps graph-level data into continuous vectors. Researchers usually utilize graph regression to formulate molecular property predictions [77]. Graph pooling has been an important topic in graph-level tasks [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], which generally utilize a hierarchical way to refine the graph structures [40, 41]. In this work, we generally study the performance of graph pooling on graph-level tasks and validate the effectiveness of graph pooling approaches in most cases.
@@ A.2 Node Classification
Node classification aims to classify each node in a graph based on its attributes and relationship with the other nodes [78, 79]. Node classification has various applications in the real world, including social network analysis [80], knowledge graphs [81], bioinformatics [82] and online commerce services [83]. Graph neural networks have been widely utilized to solve the problem by learning semantics information across nodes by neighborhood propagation. Since graph pooling would reduce the number of nodes, recent works utilize a U-Net architecture [84], which involves down-sampling and up-sampling with residue connections. In this work, we systematically evaluate the performance of graph pooling on node classification and observe that graph pooling has limited improvement compared with basic graph neural network architectures.
@@ Appendix B Details of Selected Pooling Methods
TopKPool [25]. TopKPool utilizes the attention mechanism to learn the scores of different nodes and then selects the nodes with top scores, which can learn important local portions from original graphs.
SAGPool [26]. SAGPool utilizes a different graph neural network to learn importance scores, which can guide the pooling process effectively.
ASAPool [27]. ASAPool considers the neighboring subgraphs to represent nodes and then adopts the attention mechanism to generate subgraph representations. The importance nodes are selected by a graph neural networks with local extremum information.
PANPool [28]. PANPool constructs the maximal entropy transition (MET) matrix based on graph Laplacian, which can generate importance scores for different nodes.
COPool [29]. COPool learns pooled representations from the complimentary edge and node views. The edge view comes from high-order semantics information while the node view stems from importance scores from the cut proximity matrix.
CGIPool [30]. CGIPool incorporates the mutual information optimization into graph pooling, which can enhance the graph-level relationships between the original graph and the pooled graph.
KMISPool [31]. KMISPool incorporates the Maximal k-Independent Sets (k-MIS) into graph pooling, which can detect the importance nodes in the graph with topological preserved.
GSAPool [32]. GSAPool integrates both structural and attribute information in different components. These scores from different components are then fused to guide the graph pooling process.
HGPSLPool [33]. HGPSLPool not only utilizes graph pooling to determine important nodes from the original graph, but also leverages graph structure learning to explore the topological information in the pooled graph.
AsymCheegerCutPool [34]. AsymCheegerCutPool conducts graph clustering to generate the assignment of each node according to graph total variation (GTV). Each cluster is aggregated during graph pooling in a hierarchical manner.
DiffPool [35]. DiffPool introduces a learnable soft assignment of each node during graph clustering, and then maps each cluster into the coarsened nodes in the pooling graph.
MincutPool [36]. MincutPool relaxes the classic normalized mincut problem into a continuous fashion, and then optimizes a graph neural network to achieve this. The graph clustering results are adopted to guide the graph pooling process.
DMoNPool [37]. DMoNPool introduces an objective based on modularity for graph clustering and then adds a regularization term to avoid trivial solutions during optimization. Similarly, graph clustering results are leveraged for graph pooling.
HoscPool [38]. HoscPool combines higher-order relationships in the graph with graph pooling based on motif conductance. It minimizes a relaxed motif spectral clustering objective and involves multiple motifs to learn hierarchical semantics.
JustBalancePool [39]. JustBalancePool consists of two components. On the one hand, it aims to reduces the local quadratic variation during graph clustering. On the other hand, it involves a balanced term to reduce the risk of degenerate solutions.